A woman has a son and a daughter. The mother is three times as old as her son. The son is seven years older then the daughter. Three years from now the boy will be twice as old as his sister will be. What are the three linear equations for this problem ?
Found 2 solutions by FrankM, lwsshak3:
Answer by FrankM(1040)

M=3S
S=D+7
S+3=2(D+3)
Above is the answer. To complete the problem, we can add 3 to the second equation, both sides.
S+3=D+10
Then set an equation as follows:
2(D+3)=D+10
2D+6=D+10
D=4, S=11, M=33

Answer by lwsshak3(11628)

A woman has a son and a daughter. The mother is three times as old as her son. The son is seven years older then the daughter. Three years from now the boy will be twice as old as his sister will be. What are the three linear equations for this problem ?
***
let x=son's age
3x= mother's age
x-7=daughter's age
..
3 yrs from now:
x+3=son's age
x-7+3=x-4=daughter's age
x+3=2(x-4)
x+3=2x-8
x=11
son's age=11
mother's age=33
daughter's age=4